# A Hybrid Supervised and Self-Supervised Graph Neural Network for Edge-Centric Applications

### v1

@@ I Introduction
Graphs are versatile structures used to model relationships between entities, particularly in scenarios where data does not have an Euclidean structure. Each node in a graph represents an entity, and edges describe relationships between pairs of nodes. Furthermore, nodes and edges can have features that describe properties associated with them. This flexibility makes graphs useful in various applications, including social networks [1], recommendation systems[2, 3], and bioinformatics, where they are employed to model intricate biological interactions [4].
Early approaches to graph analysis focused on using metrics such as connectivity, centrality, and other structural properties to describe graph features [5]. While useful for analyzing static properties, these methods do not learn from the graph in a way that captures complex relationships or patterns within the data. One of the earliest methods for generating node embeddings from graphs is DeepWalk [6]. This approach uses random walks to explore each node’s local neighborhood and then applies a skip-gram model to convert these walks into dense embeddings. However, it does not leverage edge weighting, which can limit its applicability. LINE [7] follows, focusing on preserving both first- and second-order proximity in the embeddings but also does not consider edge features. Node2Vec [8] extends the random walk approach by adjusting the exploration strategy for more flexible neighborhood sampling; however, it relies on fixed, non-learnable edge weights, which limits its adaptability. Despite their successes, these approaches have critical limitations: they do not share parameters across nodes, leading to inefficiencies as graphs grow, and they struggle to generalize to dynamic or unseen graphs [9]. These shortcomings underscore the need for more advanced methods capable of capturing both node and edge information in a flexible, scalable manner, setting the stage for more powerful models.
Graph Neural Networks (GNNs) have emerged as a powerful approach for addressing the limitations of traditional methods by learning from the structure of graph neighborhoods [10]. These models aim to generate new representations for entities based on their connections within the graph, leveraging the inherent relationships between nodes and edges. There are two primary approaches for constructing GNNs: spectral and spatial methods, each with its own advantages, drawbacks, and applications.
Spectral methods define graph convolutions using spectral graph theory, where the graph is transformed into the spectral domain, typically employing the graph Fourier transform [11, 12]. By leveraging the spectral components of the Laplacian matrix, these methods capture global patterns in the graph. However, they can be computationally expensive, especially for large-scale graphs, due to the need for eigenvalue decomposition of the Laplacian matrix [13, 14]. Furthermore, spectral-based models often struggle to generalize to graphs with different structures, limiting their applicability in inductive learning tasks [15, 16]. To address these issues, the GCN/ChebNet method was proposed, which uses polynomial approximations of spectral filters to reduce computational overhead [12]. However, these methods require the graph to remain unchanged during the learning process, limiting their flexibility in dynamic graph scenarios.
Spatial approaches define convolutions directly on the graph by aggregating information from neighboring nodes [10]. These methods propagate information locally, with each node passing messages to its neighbors, refining node representations iteratively [15]. Spatial methods in graph learning offer key advantages over spectral approaches, particularly in computational efficiency and flexibility. Unlike spectral methods, which require costly eigenvalue decomposition of the graph Laplacian, spatial methods operate directly on the graph structure, making them more scalable for large graphs with millions of nodes and adaptable to dynamic graphs without recalculating eigenvalues. They also generalize better to unseen nodes or graphs during training, a process known as inductive learning. This is particullarly important in scenarios where new data is constantly encountered [17]. Spatial methods also allow edge features into the learning process, which can provide relevant information for modeling complex relationships between nodes.
One of the most recent approaches in this field is the Message Passing Neural Network (MPNN) [18], which has proven effective across tasks ranging from protein-protein interaction prediction to gene regulatory network analysis [19, 20]. MPNNs are a class of GNN designed to handle graph-based tasks by iteratively passing messages between nodes and updating node representations based on the messages sent by their local neighborhood. The core idea is to aggregate information from a node’s neighbors and use this information to update the node’s feature representation. This process is particularly powerful for capturing the structural dependencies within graphs, making MPNNs highly versatile.
Formally, MPNNs operate through three main phases: message passing, node updating, and readout. At step $t$, the features $\mathbf{x}_{p}^{(t)}$ of node $p$ are updated by aggregating messages from its neighboring nodes according to:
$m_{p}^{(t+1)}=\sum_{q\in\mathcal{N}(p)}M_{t}\left(\mathbf{x}_{p}^{(t)},\mathbf{x}_{q}^{(t)},e_{pq}\right)$ (1)
where $\mathcal{N}(p)$ denotes the set of neighbors of node $p$, and $M_{t}$ is the message function that depends on both node features and edge features $e_{pq}$. The features of the central node $p$ are then updated using information from its neighboring nodes:
$\mathbf{x}_{p}^{(t+1)}=U_{t}\left(\mathbf{x}_{p}^{(t)},m_{p}^{(t+1)}\right)$ (2)
where $U_{t}$ is the node update function, which typically combines the features of neighboring nodes with aggregated messages, using schemes such as sum, mean, or maximum to update each node’s representation. After several iterations of message passing, the final node states can be aggregated using a readout function:
$\hat{y}=R\left(\{\mathbf{x}_{p}^{(K)}|p\in\mathcal{G}\}\right)$ (3)
The readout function $R$ in Eq 3 aggregates information from all nodes to produce a graph-level or sub-graph-level representation $\hat{y}$, which can be used for classification or regression tasks involving graph, nodes or edges. This architecture has proven highly effective in many applications, yet it has limitations in handling edge-centric tasks, which involve learning relationships between pairs of nodes rather than individual nodes.
Despite the success of MPNNs in node-level tasks, significant challenges remain when it comes to generalizing to large-scale or highly complex graphs and addressing edge-centric tasks such as regression and classification on graph edges. Many traditional GNN models, including MPNNs, focus primarily on node-level features while underutilizing the information contained in edge attributes, which is particularly relevant in fields like bioinformatics, where interactions between entities often occur at the edge level [4].
Based on the strengths of MPNNs, in this work we proposes a new MPNN-based model trained with a custom loss function designed to handle both supervised and self-supervised learning. Our model incorporates both node and edge features into the prediction process, making it capable of handling complex relationships between entities in graphs. Furthermore, we employ an attention mechanism inspired by transformers [21], which enhances the model’s ability to capture intricate dependencies within the graph. This architecture also proves robust even when simple one-hot encodings are used for node features, allowing the model to perform effectively even in cases where node information is sparse or missing.
@@ II-A Graph and subgraph definitions
Let $\mathcal{G}=\{v,e\}$ be a graph, where $v$ is the set of $|v|$ nodes in the network, and $e$ denotes the set of edges connecting the theme. Let $\mathbf{X}\in\mathbb{R}^{|v|\times d}$ be the feature matrix of nodes in $\mathcal{G}$, where $d$ is the number of features for each node. Furthermore, let $\mathcal{E}\in\mathbb{R}^{|e|\times q}$ denote the edge feature matrix, where each row represents the features associated with a specific edge in the graph. These edge features may or may not exist and can be a vector of arbitrary size $q$. Figure 1a shows an example of a typical graph and its elements.
@@ II-C Embedding generation
The tokenizer takes the node features $\mathbf{X}^{\prime}_{\mathcal{G}^{\prime}_{i,j}}$ as input and projects them into a more suitable space for inference using a Multi-Layer Perceptron (MLP), yielding a new representation $\widetilde{\mathbf{X}}^{\prime}_{\mathcal{G}^{\prime}_{i,j}}$. This representation is then processed by the NEAConv layer, a message-passing neural network specifically designed for edge-centric tasks. In this layer, custom message and update functions are defined to align with our model architecture.
@@
These nonlinear projections are performed by three separate perceptrons, each consisting of a linear layer followed by a sigmoid activation: The query vector $\mathbf{Q}$ is generated from the concatenation of the target node features $\widetilde{\mathbf{X}}^{\prime}_{j}$ and, when available, the edge features $\mathbf{E}^{\prime}_{\mathcal{G}^{\prime}_{i,j}}$, capturing the relevance of the target node within the context of its surrounding neighborhood. The key vector $\mathbf{K}$ is derived from the source nodes’ features $\widetilde{\mathbf{X}}^{\prime}_{\mathcal{G}^{\prime}_{i,j}}$, encoding essential information about the source nodes to be compared with the query vector. The value vector $\mathbf{V}$, also generated from the source nodes’ features $\widetilde{\mathbf{X}}^{\prime}_{\mathcal{G}^{\prime}_{i,j}}$, carries the information that will be propagated to the target node. The vector weight $\mathbf{w}$ which represents the importance of each source node’s information for the target node, is calculated as:
@@ II-D Prediction step
The prediction module in our model is a multilayer perceptron (MLP) designed to process the generated node embeddings and produce task-specific predictions. To ensure that the model’s predictions are invariant to the permutation of the nodes, we first perform a reordering of the features from the node embeddings. This reordering helps in capturing the important features across different nodes while maintaining the permutation invariance property. Initially, we extract the minimum and maximum values along the feature dimension of the node embeddings. These extracted features are then concatenated to form a new feature representation that encapsulates the critical information from the node embeddings. The architecture of the MLP consists of three linear layers, each followed by an activation function, to produce the final prediction output.
@@ II-E Loss function
where $\alpha$, $\beta$, and $\gamma$ are weighting coefficients.
@@
In addition to the supervised terms, the loss function includes a self-supervised component, $\mathcal{L}_{\text{cosine\_pred}}$, which aligns the predicted output $\hat{Y}$ with the cosine similarity of the node embeddings $\widetilde{Y}$. This self-supervised term helps the model learn robust representations even for nodes with unknown or incomplete information by structuring the embeddings in a meaningful way. As a result, the model becomes more adaptable to scenarios where data is missing, improving its generalization and performance. During training, if an input pattern does not hace true labels assigned, only the third term will be used, allowing the model to learn in a unsupervized way.
@@ III Datasets
Experimental evaluation of the proposed model was performed using three bioinformatic tasks: protein-protein interaction prediction, Gene Ontology (GO) annotation prediction, and metabolic compounds structural similarity prediction. The first task entails identifying potential interactions between proteins by analyzing structural, sequence, and functional data. The second task involves assigning an annotation of molecular functions, biological processes, or cellular components to genes or proteins based on the annotations of the known genes and all genes expression data. Finally, the last task consists of predicting the similarity between metabolic compounds that participate in the same metabolic network based on the structure of the neighborhood they share, without considering their molecular structure.
III-A Protein-Protein Interaction
Five datasets [^1] http://www.csbio.sjtu.edu.cn/bioinf/LR_PPI/Data.htm used by Yang et al. [22] were taken into account for this task. HPRD dataset involves  36k pairs of known protein interactions (positive patterns), and  36k pairs of artificial interactions constructed by randomly pairing proteins from different subcellular locations (negative patterns). Human dataset includes  9k proteins, and  72k interactions; E. coli dataset comprises  2k proteins and 13k interactions; Drosophila dataset involves  7k proteins and  44k interactions; C. elegans dataset consists of  2,5k proteins, with  8k interactions. In all cases, datasets are balanced including equal number of positive and negative interactions.
The graph for each dataset was constructed by first calculating the similarity matrix between proteins using ClustalO [23]. This similarity matrix was then used by the k-Nearest Neighbors Graph (KNNG) algorithm [24] to create an undirected graph, with the number of neighbors $k$ determined experimentally. Various $k$ values were tested, keeping the model configuration fixed, to identify the optimal $k$ for capturing meaningful connections in the graph. Node features were build from sequence representations generated by the The Composition of Triads (CT) method, proposed by Shen et al. [25], represents protein sequences by projecting them into a vector space that counts the frequency of triad types. Initially, amino acids are grouped into seven categories based on dipole and side chain volume. Each amino acid in the sequence is then replaced by its category label, transforming the protein sequence into a sequence of integers. A sliding window of size three moves along this sequence to count occurrences of each triad type, treating every three consecutive amino acids as a unit. With seven categories, this yields 343 unique triad combinations. Similarity matrix between proteins were used as edge features.
III-B Gene Ontology (GO) terms
For this task, the objective is to predict new annotations for genes with unknow labels in the GO. The model was evaluated on three organisms: Saccharomyces cerevisiae, Arabidopsis thaliana, and Dictyostelium discoideum. These organisms were selected due to their extensive functional annotations, ensuring precise performance evaluation. The datasets used in this work are described in [26].
Gene expression data for the YEAST dataset, involving Saccharomyces cerevisiae, was sourced from [27]. From an initial 2,467 genes, 605 with 79 microarray values were selected after excluding those with missing data. GO annotations were applied following CAFA challenge guidelines [28]. After filtering out terms annotated to fewer than three genes, 422 genes remained for Biological Process (BP), 386 for Molecular Function (MF), and 442 for Cellular Component (CC). The ARA dataset, involving circadian-regulated genes in Arabidopsis thaliana under cold stress, comprised 1,546 genes, with 656 retaining BP annotations after filtering. Ultimately, 521 genes remained for BP, 315 for MF, and 740 for CC [29]. Gene expression data from Dictyostelium discoideum over a 24-hour development cycle constituted the DICTY dataset [30]. Initially, 707 genes with BP annotations were selected, and after filtering, 652 genes remained for BP, 366 for MF, and 630 for CC.
The expression vectors of each gene were used as node features, representing the gene’s expression profile, while the semantic similarity between each gene pair was used as edge features and to construct the $\mathcal{G}$ graph using the KNNG, which was then transformed into an undirected graph. Various $k$ values were tested, keeping the model configuration fixed, to identify the optimal $k$. The model is used to predict the semantic similarity for all unknown genes.
@@ III-C Metabolic pathways
For this study, the objective is to predict the similarity of compounds included in metabolic pathways. We expanded upon our previous work [31] on the glycolysis metabolic pathway to include five additional metabolic pathways. Now encompasses the following pathways: Glycolysis, Starch and sucrose metabolism, Pentose phosphate pathway, Citrate cycle (TCA cycle), Pyruvate metabolism and Propanoate metabolism.
This expanded dataset comprises a total of 207 compounds, of which 174 have known structures and 33 have unknown structures. The dataset was constructed following the methodology in our previous work [31, 32]. Data were extracted from the KEGG [^2] https://www.genome.jp/kegg/ (v95.2) database. For compounds with known structures, molecular structures were downloaded in SMILES format from the PubChem [^3] https://pubchem.ncbi.nlm.nih.gov/ database. The RDKit [^4] https://www.rdkit.org library was used to calculate MACC keys [33] fingerprints for all compounds with known structures, as demonstrated in our previous work [32], where we showed the effectiveness of these fingerprints for this task. The Tanimoto coefficient [34] was then applied to calculate the similarity between compounds, as defined by:
$T(\mathbf{c}_{i},\mathbf{c}_{j})=\frac{\sum_{k}(\mathbf{c}_{i}^{k}\wedge\mathbf{c}_{j}^{k})}{\sum_{k}(\mathbf{c}_{i}^{k}\vee\mathbf{c}_{j}^{k})}$ (9)
where $\wedge$ and $\vee$ are the binary operators $and$ and $or$, respectively, and $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ are binary representations of the structures of compounds $i$ and $j$. The Tanimoto coefficient takes values in the range $[0,1]$ and calculates the proportion of shared features between the two structures. In total, 21,528 patterns were defined, resulting from the pairwise combinatorics of the 174 compounds with known structures and their interactions with compounds of unknown structure. Of these patterns, 5,390 are related to compounds with unknown structures.
For this dataset, one-hot vectors were used as node features to represent the compounds, while a one-hot vector based on the enzyme family that mediates the reaction served as edge features. These enzyme families are determined by the first digit of their KEGG EC number, ranging from 1 to 7 (e.g., 1 for Oxidoreductases, 2 for Transferases, and so forth). The EC numbers, standardized by the IUBMB/IUPAC, classify enzymes based on their biochemical reactions and are used in KEGG as ortholog identifiers (KOs) to link reactions with protein sequence data. In cases where multiple enzymes were involved, their one-hot vectors were aggregated. The construction of the metabolic pathway graph involved representing each compound as a node and connecting two nodes with an edge if they participated in a reaction as substrate and product. A comprehensive graph was built by integrating all reactions from the six specified metabolic pathways.
@@ IV Results
For GO terms prediction results are reported according to the CAFA rules, with the maximum F1-measure ($\text{F}1_{max}$), which considers predictions across the full spectrum from high to low sensitivity. The calculation of $\text{F}1_{max}$ is conducted through a systematic approach to evaluate the quality of predictions in biological functions. First, the results of the predictions and the actual annotations for the functions being assessed are collected. Next, precision and recall are calculated for each function, considering all possible thresholds to observe how these metrics vary. The decision thresholds are then adjusted, and for each threshold, the F1 measure, which combines precision and recall into a single metric, is computed. $\text{F}1_{max}$ is defined as the maximum F1 value obtained by varying the thresholds, allowing for the identification of the threshold that maximizes this metric. Ultimately, a higher $\text{F}1_{max}$ value indicates a better balance between precision and recall, reflecting greater effectiveness in the predictions made.

### v2

@@ I Introduction
Graphs are versatile structures used to model relationships between entities in non-Euclidean domains. They underpin applications ranging from social networks [1] and recommendation systems [2, 3] to bioinformatics [4]. Early shallow embedding methods such as DeepWalk [5], LINE [6], and Node2Vec [7] demonstrated the value of data-driven representations but share two well-known limitations: parameter redundancy and limited generalization [8]. Graph Neural Networks (GNNs) overcome these drawbacks by parameter sharing and inductive learning [9]. Within GNNs, spectral and spatial formulations provide complementary views; the reader is referred to [10, 11, 12, 13, 14, 15] for canonical treatments.
Message Passing Neural Networks (MPNNs) generalize earlier GNN variants by formalizing graph convolutions as iterative message-passing steps [16]. At iteration $t$, node $p$ aggregates information from its neighbours $\mathcal{N}(p)$ via
$\begin{split}m_{p}^{(t+1)}&=\sum_{q\in\mathcal{N}(p)}M_{t}\!\bigl{(}\mathbf{x}_{p}^{(t)},\mathbf{x}_{q}^{(t)},e_{pq}\bigr{)},\\ \mathbf{x}_{p}^{(t+1)}&=U_{t}\!\bigl{(}\mathbf{x}_{p}^{(t)},m_{p}^{(t+1)}\bigr{)},\end{split}$ (1)
where $\mathbf{x}_{p}^{(t)}$ and $\mathbf{x}_{q}^{(t)}$ are the node features at step $t$, $e_{pq}$ is an optional edge feature, $M_{t}$ is the message function, $m_{p}^{(t+1)}$ is the aggregated message, and $U_{t}$ is the update function. After $T$ iterations, an order-invariant readout $R\!\bigl{(}\mathbf{x}_{p}^{(T)}\bigr{)}$ (e.g., sum or mean pooling) produces the final prediction at node or graph level. Equation (1) establishes the notation used throughout the paper.
Motivation and contribution.
Existing MPNNs excel at node-level tasks but still under-utilize edge attributes and require large labelled datasets for supervision. Edge-centric problems—typical in bioinformatics, where interactions are often encoded on edges [4]—remain challenging. This work addresses these gaps by:
  1. 1.
    Proposing an MPNN-based architecture that natively treats node and edge features with a shared attention mechanism inspired by transformers [17];
  2. 2.
    Introducing a hybrid loss that couples supervised and self-supervised objectives, enabling learning even when node features are sparse (e.g. one-hot encodings);
  3. 3.
    Demonstrating competitive performance on tasks that require edge-level regression and classification, including protein–protein interaction prediction.
The remainder of this paper is organised as follows. Section II details the materials and methods that ground the proposed model; Section III introduces the benchmark datasets employed in our experiments; Section IV reports and discusses the empirical results and ablation studies; finally, Section V summarises the main contributions and outlines future research directions.
@@ II-A Graph and subgraph definitions
Let $\mathcal{G}=\{v,e\}$ be a graph, where $v$ is the set of $|v|$ nodes in the network, and $e$ denotes the set of edges connecting them. Let $\mathbf{X}\in\mathbb{R}^{|v|\times d}$ be the feature matrix of nodes in $\mathcal{G}$, where $d$ is the number of features for each node. Furthermore, let $\mathcal{E}\in\mathbb{R}^{|e|\times q}$ denote the edge feature matrix, where each row represents the features associated with a specific edge in the graph. These edge features may or may not exist and can be a vector of arbitrary size $q$. Figure 1a shows an example of a typical graph and its elements.
@@ II-C Embedding generation
The tokenizer takes the node features $\mathbf{X}^{\prime}_{\mathcal{G}^{\prime}_{i,j}}$ as input and projects them into a more suitable space for inference using a Multi-Layer Perceptron (MLP), yielding a new representation $\widetilde{\mathbf{X}}^{\prime}_{\mathcal{G}^{\prime}_{i,j}}$. This representation is then processed by a single NEAConv layer, a message-passing neural network specifically designed for edge-centric tasks. In this layer, custom message and update functions are defined to align with our model architecture.
@@
These nonlinear projections are performed by three separate perceptrons, each consisting of a linear layer followed by a sigmoid activation. Because the propagated information comes exclusively from first-order neighbours rather than from an ordered sequence, the incoming messages are intrinsically unsorted; therefore, no positional encoding is applied. The query vector $\mathbf{Q}$ is generated from the concatenation of the target node features $\widetilde{\mathbf{X}}^{\prime}_{j}$ and, when available, the edge features $\mathbf{E}^{\prime}_{\mathcal{G}^{\prime}_{i,j}}$, capturing the relevance of the target node within the context of its surrounding neighborhood. The key vector $\mathbf{K}$ is derived from the source nodes’ features $\widetilde{\mathbf{X}}^{\prime}_{\mathcal{G}^{\prime}_{i,j}}$, encoding essential information about the source nodes to be compared with the query vector. The value vector $\mathbf{V}$, also generated from the source nodes’ features $\widetilde{\mathbf{X}}^{\prime}_{\mathcal{G}^{\prime}_{i,j}}$, carries the information that will be propagated to the target node. The vector weight $\mathbf{w}$ which represents the importance of each source node’s information for the target node, is calculated as:
@@ II-D Prediction step
The prediction module in our model is a multilayer perceptron (MLP) designed to process the generated node embeddings and produce task-specific predictions. To ensure that the model predictions are invariant to the permutation of the nodes, we first perform a reordering of the features from the node embeddings. This reordering helps in capturing the important features across different nodes while maintaining the permutation invariance property. Initially, we extract the minimum and maximum values along the feature dimension of the node embeddings. These extracted features are then concatenated to form a new feature representation that encapsulates the critical information from the node embeddings. The architecture of the MLP consists of three linear layers, each followed by an activation function, to produce the final prediction output.
@@ II-E Loss function
where $\alpha$, $\beta$, and $\gamma$ denote weighting coefficients. Preliminary experiments showed that the three loss components are of the same order of magnitude; accordingly, we fixed $\alpha=\beta=\gamma=1$ to avoid privileging any single term. These coefficients may, however, be treated as hyper-parameters and tuned for particular datasets in future work.
@@
In addition to the supervised terms, the loss function includes a self-supervised component, $\mathcal{L}_{\text{cosine\_pred}}$, which aligns the predicted output $\hat{Y}$ with the cosine similarity of the node embeddings $\widetilde{Y}$. This self-supervised term helps the model learn robust representations even for nodes with unknown or incomplete information by structuring the embeddings in a meaningful way. As a result, the model becomes more adaptable to scenarios where data is missing, improving its generalization and performance. During training, if an input pattern does not have true labels assigned, only the third term will be used, allowing the model to learn in a unsupervised way.
@@ III Datasets
Experimental validation covers three widely used bioinformatic tasks in the literature: Protein–Protein Interaction (PPI) prediction, Gene Ontology (GO) annotation, and metabolic-compound similarity.
III-A Protein–Protein Interaction
Five balanced PPI benchmarks introduced by Yang et al. [18] [^1] http://www.csbio.sjtu.edu.cn/bioinf/LR_PPI/Data.htm are employed: HPRD, Human, E. coli, Drosophila and C. elegans. For every dataset a $k$-nearest-neighbour graph (KNNG) is built from ClustalO sequence similarities [19, 20]; $k$ is tuned on the validation split. Node features use Composition-of-Triads (343-dim.) representations [21], and pairwise similarity scores act as edge attributes.
III-B Gene Ontology Terms
Following the EXP2GO protocol [22], GO annotation transfer is evaluated on Saccharomyces cerevisiae, Arabidopsis thaliana and Dictyostelium discoideum. Expression profiles (79–740 measurements per gene) from  [23, 24, 25] constitute node features, while semantic similarities define weighted edges. Annotations are filtered as in CAFA [26]. The KNNG construction and $k$ selection mirror the PPI pipeline.
@@ III-C Metabolic Pathways
Expanding our previous glycolysis study [27, 28], the benchmark now encompasses six KEGG pathways: Glycolysis, Starch and sucrose metabolism, Pentose phosphate pathway, Citrate cycle (TCA cycle), Pyruvate metabolism, and Propanoate metabolism. The resulting graph contains 207 compounds, of which 174 have known SMILES structures and 33 lack structural information (unknown compounds). MACCS fingerprints [29] computed with RDKit [^2] https://www.rdkit.org serve as node features for the known compounds; one-hot placeholders mark unknown structures. Edge weights are Tanimoto coefficients [30], and enzyme-family one-hot vectors (EC 1–7) supply additional edge attributes. In total, 21 528 compound pairs are generated, including 5 390 pairs that involve at least one unknown compound.
@@ IV Results
For GO terms prediction results are reported according to the CAFA rules [26], with the maximum F1-measure ($\text{F}1_{max}$), which considers predictions across the full spectrum from high to low sensitivity. The calculation of $\text{F}1_{max}$ is conducted through a systematic approach to evaluate the quality of predictions in biological functions. First, the results of the predictions and the actual annotations for the functions being assessed are collected. Next, precision and recall are calculated for each function, considering all possible thresholds to observe how these metrics vary. The decision thresholds are then adjusted, and for each threshold, the F1 measure, which combines precision and recall into a single metric, is computed. $\text{F}1_{max}$ is defined as the maximum F1 value obtained by varying the thresholds, allowing for the identification of the threshold that maximizes this metric. Ultimately, a higher $\text{F}1_{max}$ value indicates a better balance between precision and recall, reflecting greater effectiveness in the predictions made.
@@
Figure 4: Comparison of our model with SVGAE, Siamese Residual RCNN, and EnAmDNN on five PPI organisms. Each boxplot represents the distribution of the F1 scores obtained from the 10 times 5-fold cross-validation. The boxplots show the mean F1 score for each k-fold.